# XTRUST: On the Multilingual Trustworthiness of Large Language Models

### v1

@@ 4.9 Evaluation on Misinformation
Baichuan 1.6 14.9 0 4.1 0 12.0 4.9 8.2 2.9 6.2 5.48
Gemini 28.6 42.3 20.3 51.8 1.2 78.8 0 34.4 0 37.3 29.47
Davinci 1.6 0.8 0.8 0 0 0 0 0 1.2 0.4 0.48
ChatGPT 9.5 29.8 26.1 8.7 0 47.3 7.5 22.8 17 28.6 19.73
GPT-4 49.3 12.4 51.4 50.6 0.4 90 16.5 62.6 2.0 48.5 38.37
@@ 4.10 Evaluation on Privacy
Baichuan 17.67 27.95 26.29 24.02 19.59 25.62 18.48 21.75 22.67
Gemini 29.03 41.37 44.83 50.23 48.6 57.4 58.16 51.02 47.58
Davinci 34.58 34.16 33.93 38.34 34.52 31.42 28.75 31.54 33.41
ChatGPT 78.34 81.65 78.35 79.71 73.92 82.18 74.83 77.71 78.34
GPT-4 80.6 85.02 83.01 85.32 81.68 85.22 83.72 84.27 83.61
@@
Results. For the PII detection task, as shown in Table 10, Baichuan, Gemini, and Davinci demonstrate weak performance, leaving significant room for improvement in preventing data leakage during conversations. In contrast, ChatGPT and GPT-4 exhibit strong performance in protecting PII during interactions. Notably, ChatGPT performs best at safeguarding phone numbers, though it still shows vulnerabilities by leaking sensitive information such as addresses and email addresses. For the privacy event detection task, GPT-4 outperforms all other LLMs across all privacy events, surpassing the second-best model, ChatGPT, by a margin of 7.09%. In summary, GPT-4 excels at protecting private information, demonstrating its robustness and superior ability to detect and handle inappropriate instructions.
@@ 4.11 Evaluation on Machine Ethics
Baichuan 0-shot_ETHICS 0 65.6 55.2 64.1 0 55.7 63.6 59.7 62.6 58.7 48.52
5-shot_ETHICS 0 69.1 60.1 62.1 0.9 41.7 58.7 62.6 54.7 56.2 46.61
@@
5-shot_JC 0 44.7 12.4 36.3 0 21.3 48.2 49.2 15.9 42.7 27.07
Gemini 0-shot_ETHICS 12.9 60.1 8.9 6.9 0.9 7.9 8.4 5.4 8.9 9.9 13.02
5-shot_ETHICS 0 41.7 3.9 4.4 6.4 5.4 1.4 7.9 3.9 7.4 8.24
0-shot_JC 7.4 32.8 3.9 2.4 0 2.4 4.4 4.4 5.4 3.4 6.65
5-shot_JC 0.9 32.84 4.4 3.9 4.4 1.4 2.9 4.4 10.9 2.4 6.844
@@
5-shot_ETHICS 0 0.9 0 4.4 0 2.9 0 6.9 10.9 3.9 2.99
ChatGPT 0-shot_ETHICS 0 66.1 69.1 66.6 4.9 71.6 63.1 72.1 62.1 71.1 54.67
5-shot_ETHICS 8.4 62.1 67.1 68.6 4.9 69.1 63.6 69.6 58.7 74.1 54.62
@@
5-shot_JC 5.4 53.2 49.2 50.2 26.8 50.2 52.2 48.7 61.1 54.7 45.17
GPT-4 0-shot_ETHICS 0 81.5 76.1 77.1 0.4 77.1 71.1 79.1 76.1 79.6 61.81
5-shot_ETHICS 0 81.0 78.1 80.0 0.4 77.6 71.6 77.1 77.1 78.6 62.15
@@
Setup. To assess how well LLMs can distinguish between moral and immoral behaviors, we compare their moral recognition performance using two morality datasets: ETHICSGui et al. (2021) and Jiminy Cricket (JC)Hariri (2023). The ETHICS dataset evaluates a model’s understanding of fundamental ethical concepts, including justice, virtue, deontology, utilitarianism, and commonsense morality. It comprises five subsets, each focusing on different normative factors. For this study, we follow the approach in DecodingTrust and select the commonsense morality subset for evaluation. In the few-shot setting, we provide five training samples as demonstrations, each consisting of a commonsense morality scenario and its label. The LLM is required to classify each sample as either "wrong" or "not wrong," with classification accuracy serving as the evaluation metric. The JC dataset is designed to assess an agent’s ability to act morally in text-based games, featuring more diverse scenarios and annotations. Each sample includes a scenario from a text-based game, accompanied by a threefold label: (1) the moral valence of the action (good, bad, or neutral), (2) the focal point (whether the action benefits or harms the agent or others), and (3) the ordinal degree (a ranking of how good or bad the action is on a scale of 1 to 3). In the few-shot setting, three training samples are provided as demonstrations. The LLMs are tasked with determining the moral valence of each sample (good, bad, or neutral), and classification accuracy is used as the evaluation metric. For machine ethics, we evaluate the LLMs in both zero-shot and five-shot settings.
@@ 5 Conclusion
In this study, our primary focus is on exploring the multilingual trustworthiness capabilities of LLMs. However, two key limitations prevent us from providing a comprehensive assessment of LLMs’ trustworthiness in practical applications. First, although we evaluate five widely-used LLMs, we were unable to assess other open-source and proprietary models, such as Llama and Claude, which also support multiple languages. Second, our evaluation does not encompass all non-English languages. We plan to expand our evaluation in future iterations of this study, with the support of the research community.
Appendix A Prompt design
Figure 3: An example of the input prompt.
@@ Appendix A Prompt design
In most instances, LLMs adhere to the provided instructions and generate expected responses. However, as noted in previous studies Zhou et al. (2023); Huang et al. (2023), LLMs may occasionally refuse to answer or produce hallucinations unrelated to the input. In this study, such responses are considered incorrect for the purposes of evaluation, particularly in classification tasks.
Appendix B Evaluated Models
@@
GPT-3.5-turbo Undisclosed api 1106 Multi-Lans OpenAI
@@
Appendix C Demonstration of the illegal activity evaluation for LLMs
In Figure 4, we present an example of illegal activity evaluation for LLMs.
Figure 4: An example of illegal activity evaluation.
Appendix D Demonstration of the hallucination evaluation for LLMs
We have presented an example of the hallucination evaluation for LLMs in Figure 5.
Figure 5: An example of hallucination evaluation.
Appendix E Demonstration of the OOD evaluation for LLMs
In Figure 6, we present an example of the OOD evaluation for LLMs.
Figure 6: An example of OOD evaluation.
Appendix F Demonstration of the mental health evaluation for LLMs
In Figure 7, we present an example of the mental health evaluation for LLMs.
Figure 7: An example of the mental health evaluation.
Appendix G Demonstration of the physical health evaluation for LLMs
We have included an example related to the physical health evaluation for LLMs in Figure 8.
Figure 8: An example of the physical health evaluation.
Appendix H Demonstration of the fairness and bias evaluation for LLMs
In Figure 9, we present an example of fairness and bias evaluation for LLMs. Furthermore, Figure 10 provides a visualized analysis of different prompts across diverse languages and models, showing that GPT-4 and ChatGPT surpass other language models in fairness and bias evaluation for all tested languages across different prompt types.
Figure 9: An example of the fairness and bias evaluation.
Figure 10: The fairness and bias evaluation for LLMs across three prompt types.
Appendix I Demonstration of the toxicity evaluation for LLMs
We have included an example related to the toxicity evaluation for LLMs in Figure 11.
Figure 11: An example of the toxicity evaluation.
Appendix J Demonstration of the misinformation evaluation for LLMs
We have included an example related to the misinformation evaluation for LLMs in Figure 12.
Figure 12: An example of the misinformation evaluation.
Appendix K Demonstration of the PII detection for LLMs
We have included an example of the PII detection for LLMs in Figure 13.
Figure 13: An example of the PII detection.
Appendix L Demonstration of the PrivacyEvents detection for LLMs
We present an example of the PrivacyEvents detection for LLMs in Figure 14. In Figure 15161718, and 19, we provide the analyses of privacy-related word leakage for PII and PrivacyEvents detection experiments. These results reveal that for powerful LLMs, such as GPT-4 and ChatGPT, can detect most privacy-related words for most languages. GPT-4 performs poorly in Russian and ChatGPT performs poorly in French. These results provide more insights into understanding different privacy-related words for different languages.
Figure 14: An example of the PrivacyEvents detection.
Figure 15: Analysis of privacy-related word leakage of GPT-4 for privacy detection.
Figure 16: Analysis of privacy-related word leakage of ChatGPT for privacy detection.
Figure 17: Analysis of privacy-related word leakage of Baichuan for privacy detection.
Figure 18: Analysis of privacy-related word leakage of Gemini for privacy detection.
Figure 19: Analysis of privacy-related word leakage of Davinci for privacy detection.
Appendix M Demonstration of the machine ethics evaluation for LLMs
We present an example of the machine ethics evaluation task in zero-shot setting in Figure 20.
Figure 20: An example of the machine ethics evaluation.

### v2

@@ 4.9 Evaluation on Misinformation
Baichuan 1.6 14.9 0 4.1 0 12.0 4.9 8.2 2.9 6.2 5.5
Gemini 28.6 42.3 20.3 51.8 1.2 78.8 0 34.4 0 37.3 29.5
Davinci 1.6 0.8 0.8 0 0 0 0 0 1.2 0.4 0.5
ChatGPT 9.5 29.8 26.1 8.7 0 47.3 7.5 22.8 17 28.6 19.7
GPT-4 49.3 12.4 51.4 50.6 0.4 90.0 16.5 62.6 2.0 48.5 38.4
@@ 4.10 Evaluation on Privacy
Baichuan 17.7 28.0 26.3 24.0 19.6 25.6 18.5 21.8 22.7
Gemini 29.0 41.4 44.8 50.2 48.6 57.4 58.2 51.0 47.6
Davinci 34.6 34.2 33.9 38.3 34.5 31.4 28.8 31.5 33.4
ChatGPT 78.3 81.7 78.4 79.7 73.9 82.2 74.8 77.7 78.3
GPT-4 80.6 85.0 83.0 85.3 81.7 85.2 83.7 84.3 83.6
@@
Results. For the PII detection task, as shown in Table 10, Baichuan, Gemini, and Davinci demonstrate weak performance, leaving significant room for improvement in preventing data leakage during conversations. In contrast, ChatGPT and GPT-4 exhibit strong performance in protecting PII during interactions. Notably, ChatGPT performs best at safeguarding phone numbers, though it still shows vulnerabilities by leaking sensitive information such as addresses and email addresses. For the privacy event detection task, GPT-4 outperforms all other LLMs across all privacy events, surpassing the second-best model, ChatGPT, by a margin of 7.09%. In summary, GPT-4 excels at protecting private information, demonstrating its robustness and superior ability to detect and handle inappropriate instructions. While ChatGPT performs well in certain areas, particularly phone number protection, there is still room for improvement in safeguarding all types of sensitive data across different tasks.
@@ 4.11 Evaluation on Machine Ethics
Baichuan 0-shot_ETHICS 0 65.6 55.2 64.1 0 55.7 63.6 59.7 62.6 58.7 48.5
5-shot_ETHICS 0 69.1 60.1 62.1 0.9 41.7 58.7 62.6 54.7 56.2 46.6
@@
5-shot_JC 0 44.7 12.4 36.3 0 21.3 48.2 49.2 15.9 42.7 27.1
Gemini 0-shot_ETHICS 12.9 60.1 8.9 6.9 0.9 7.9 8.4 5.4 8.9 9.9 13.0
5-shot_ETHICS 0 41.7 3.9 4.4 6.4 5.4 1.4 7.9 3.9 7.4 8.2
0-shot_JC 7.4 32.8 3.9 2.4 0 2.4 4.4 4.4 5.4 3.4 6.7
5-shot_JC 0.9 32.8 4.4 3.9 4.4 1.4 2.9 4.4 10.9 2.4 6.8
@@
5-shot_ETHICS 0 0.9 0 4.4 0 2.9 0 6.9 10.9 3.9 3.0
ChatGPT 0-shot_ETHICS 0 66.1 69.1 66.6 4.9 71.6 63.1 72.1 62.1 71.1 54.7
5-shot_ETHICS 8.4 62.1 67.1 68.6 4.9 69.1 63.6 69.6 58.7 74.1 54.6
@@
5-shot_JC 5.4 53.2 49.2 50.2 26.8 50.2 52.2 48.7 61.1 54.7 45.2
GPT-4 0-shot_ETHICS 0 81.5 76.1 77.1 0.4 77.1 71.1 79.1 76.1 79.6 61.8
5-shot_ETHICS 0 81.0 78.1 80.0 0.4 77.6 71.6 77.1 77.1 78.6 62.2
@@
Setup. To assess how well LLMs can distinguish between moral and immoral behaviors, we compare their moral recognition performance using two morality datasets: ETHICS Gui et al. (2021) and Jiminy Cricket (JC) Hariri (2023). The ETHICS dataset evaluates a model’s understanding of fundamental ethical concepts, including justice, virtue, deontology, utilitarianism, and commonsense morality. It comprises five subsets, each focusing on different normative factors. For this study, we follow the approach in DecodingTrust and select the commonsense morality subset for evaluation. In the few-shot setting, we provide five training samples as demonstrations, each consisting of a commonsense morality scenario and its label. The LLM is required to classify each sample as either "wrong" or "not wrong," with classification accuracy serving as the evaluation metric. The JC dataset is designed to assess an agent’s ability to act morally in text-based games, featuring more diverse scenarios and annotations. Each sample includes a scenario from a text-based game, accompanied by a threefold label: (1) the moral valence of the action (good, bad, or neutral), (2) the focal point (whether the action benefits or harms the agent or others), and (3) the ordinal degree (a ranking of how good or bad the action is on a scale of 1 to 3). In the few-shot setting, five training samples are provided as demonstrations. The LLMs are tasked with determining the moral valence of each sample (good, bad, or neutral), and classification accuracy is used as the evaluation metric. For machine ethics, we evaluate the LLMs in both zero-shot and five-shot settings.
@@ 5 Conclusion
In this study, our primary focus is on exploring the multilingual trustworthiness capabilities of LLMs. However, three key limitations prevent us from providing a comprehensive assessment of LLMs’ trustworthiness in practical applications. First, although we evaluated five widely-used LLMs, we were unable to assess all possible open-source and proprietary models, such as Llama and Claude, which also support multiple languages. Second, our evaluation does not encompass all available non-English languages. Third, due to significant cross-linguistic variations, we found that no single prompt engineering approach could comprehensively enhance LLM performance on XTrust across different languages. These limitations highlight the need for continued exploration in future iterations of this study to address linguistic diversity and model coverage challenges.
Ethics Statement
This study systematically evaluates LLMs across 10 multilingual trustworthiness dimensions—including legal compliance, ethical reasoning, and privacy sensitivity—using carefully curated non-sensitive datasets. To ensure ethical rigor, our methodology adheres to international AI ethics standards and explicitly prohibits models from endorsing harmful actions. Evaluations prioritize two objectives: (1) identifying systemic risks in LLM decision-making through scenario-based assessments, and (2) establishing accountability via transparent reporting of aggregated results. All resources, including multilingual evaluation frameworks and multilingual datasets, are available in a publicly accessible repository to promote reproducibility. Furthermore, we emphasize pretraining strategies optimized for multilingual generalization, aiming to advance equitable AI deployment while mitigating cross-cultural biases. This work contributes to the global discourse on AI safety by balancing technical innovation with stringent ethical safeguards.
Appendix A Appendix
A.1 Data Collection
For Illegal Activity, we randomly sampled 200 instances from the Chinese AI and Law challenge dataset (CAIL2018) Xiao et al. (2018), a large compilation of legal case records provided by the Supreme People’s Court of China (While based on Chinese law, the task focuses on cross-linguistic performance, leveraging universal legal concepts like punishing murder, and aligns with the importance of multilingual scenarios in NLP).
For Hallucination, we selected 300 samples from the Question Answering (QA) task in HaluevalLi et al. (2023), a dataset containing generated and human-annotated hallucinated examples to evaluate LLMs’ ability to detect hallucinations.
For OOD Robustness, we gathered entertainment, education, and international news from China News Service, specifically collecting articles published after September 2023 to ensure the information was not used in LLM training. A total of 200 articles were used for this evaluation.
For Mental Health, we randomly selected 200 instances from the Mental Health subset of SafetyBench Zhang et al. (2023c), which includes safety-related multiple-choice questions.
For Physical Health, we randomly selected 200 samples from SafeText Levy et al. (2022), a dataset containing 367 human-written real-life scenarios with safe and unsafe suggestions for each scenario.
For Toxicity, we randomly selected 200 instances from the toxic task prompts subset of REALTOXICITYPROMPT Gehman et al. (2020) to evaluate the toxicity generation of LLMs.
For Fairness & Bias, we selected prompts from the Stereotype Bias subset of DecodingTrust Wang et al. (2023a), asking LLMs whether they agree or disagree with the statements presented in the prompts. DecodingTrust spans 24 demographic groups (e.g., "Homosexuals," "Men," "White people") and 16 bias topics (e.g., "Sexually Transmitted Diseases," "Occupations," "Intelligence"). We selected 384 prompts (24 demographic groups × 16 bias topics) to comprehensively evaluate LLMs’ performance on fairness and bias.
For Misinformation, we used 155 questions from the Misinformation Harms subset of the Do-Not-Answer dataset Wang et al. (2023b), a taxonomy for assessing the risk of harm posed by LLMs. Additionally, we collected 86 fake news-related questions from the Internet, using a total of 241 questions in this evaluation.
For Privacy, we selected 34 data entries from the Privacy subset of DecodingTrust Wang et al. (2023a) to evaluate LLMs’ capability to prevent data leakage.
For Machine Ethics, we randomly selected 200 short samples from the test set of ETHICS Gui et al. (2021), which assesses a model’s knowledge of basic ethical concepts. We also selected 200 instances from the Jiminy Cricket dataset Hariri (2023), which evaluates an agent’s ability to make moral decisions in text-based games.
A.2 Prompt design
@@
In most instances, LLMs adhere to the provided instructions and generate expected responses. However, as noted in previous studies Huang et al. (2023), LLMs may occasionally refuse to answer or produce hallucinations unrelated to the input. In this study, such responses are considered incorrect for the purposes of evaluation, particularly in classification tasks.
Figure 4: Example input prompt for the evaluation task
A.3 Evaluated Models
@@
ChatGPT-turbo Undisclosed api 1106 Multi-Lans OpenAI
@@
A.4 Demonstrations of Evaluation Tasks
We provide visual demonstrations for each evaluation task:
1. Figure 5 shows the illegal activity evaluation example 2. Figure 6 demonstrates the hallucination evaluation 3. Figure 7 presents the out-of-distribution (OOD) evaluation 4. Figure 8 illustrates the mental health evaluation 5. Figure 9 displays the physical health evaluation 6. Figures 10 and 11 show fairness and bias evaluation examples and cross-linguistic analysis 7. Figure 12 demonstrates the toxicity evaluation 8. Figure 13 presents the misinformation evaluation 9. Figure 14 shows personal identifiable information (PII) detection 10. Figures 15 through 20 present privacy events detection and analyses across models 11. Figure 21 demonstrates the machine ethics evaluation
Figure 5: Illegal activity evaluation example.
Figure 6: Hallucination evaluation example.
Figure 7: Out-of-distribution evaluation example.
Figure 8: Mental health evaluation example.
Figure 9: Physical health evaluation example.
Figure 10: Fairness and bias evaluation example.
Figure 11: Cross-model fairness and bias analysis.
Figure 12: Toxicity evaluation example.
Figure 13: Misinformation evaluation example.
Figure 14: PII detection example.
Figure 15: Privacy events detection example.
Figure 16: GPT-4 privacy word leakage analysis.
Figure 17: ChatGPT privacy word leakage analysis.
Figure 18: Baichuan privacy word leakage analysis.
Figure 19: Gemini privacy word leakage analysis.
Figure 20: Davinci privacy word leakage analysis.
Figure 21: Machine ethics evaluation example.